OpenStudy (anonymous):

factor the polynomial and check your answer 4m^4 - 2m^2 + 8m

OpenStudy (shadowfiend):

Is there a common power of m in all three of those terms?

OpenStudy (anonymous):

nope

OpenStudy (shadowfiend):

There is. Remember that \(m^4 = m \cdot m \cdot m \cdot m\) and \(m^2 = m \cdot m\).

OpenStudy (anonymous):

what?

OpenStudy (shadowfiend):

\[4m^4 - 2m^2 + 8m = 4mmmm - 2mm + 8m\] So you can pull a single m out of all three terms: \[m(4m^3 - 2m + 8)\] There's one more common factor, it's a number though. Is there a factor common to 4, 2, and 8?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

?